OpenStudy (leahmarieed):

Choose the equation below that represents the line passing through the point (−3, −1) with a slope of 4.

Mehek (mehek14):

Equations?

OpenStudy (leahmarieed):

y = 4x − 11 y = 4x + 11 y = 4x + 7 y = 4x – 7

Mehek (mehek14):

All of the equations have a slope of 4 `4`x We need to find the y-intercept Plug the ordered pair into the equation -1 = 4 * -1 + b

Mehek (mehek14):

`-1 = 4 * -1 + b` My bad: -1 = 4 * -3 + b